shiversmekelah441:

What is the value of x in the equation StartFraction 4 Over 7 EndFraction (StartFraction 21 Over 8 EndFraction x + one-half) = negative 2 (StartFraction 1 Over 7 EndFraction minus StartFraction 5 Over 28 EndFraction x)? Write your solution in decimal form.

Vocaloid:

(4/7) * (21/8x + 1/2) = -2(1/7 - 5/28x) \[\frac{ 4 }{ 7 }(\frac{ 21 }{ 8 }x+\frac{ 1 }{ 2 })=-2(\frac{ 1 }{ 7 }-\frac{ 5 }{ 28 }x)\]